OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2-6x+8=y The question options are: (–3,– 6) B. (3,–1) C. (–3,1) D. (–1, 3) I picked c but im not sure if its correct.

OpenStudy (dumbcow):

\[y = (x^{2}-6x) +8\] \[y = (x^{2}-6x+9)+8-9\] \[y = (x-3)^{2} -1\] vertex: (3,-1)

OpenStudy (anonymous):

SO i wasnt right?

OpenStudy (dumbcow):

nope, answer is B

OpenStudy (anonymous):

Ok thanks lolz